OpenStudy (anonymous):

Solve using the guess-and-test strategy. The math club has 57 members who are in sixth or seventh grades. There are twice as many sixth graders as seventh graders. How many sixth graders are in the club? A. 38 B. 37 C. 19 D. 18

OpenStudy (perl):

Let x = # of sixth graders y = # of seventh graders x + y = 57 x = 2*y

OpenStudy (perl):

sorry I'm not familiar with guess and test strategy. use the multiple choice?

OpenStudy (kamibug):

@perl is right. :) Use an equation. Since x = 2y, you can substitute in that value and the equation would be: 2y +y = 57 Then solve for y: 3y = 57 y = 19 But remember "y" is the number of 7th graders. So now that we know what "y" is, we can plug that value into the original equation and solve for x (the # of 6th graders). x + 19 = 57 x = 38

OpenStudy (anonymous):

is it A?

OpenStudy (kamibug):

Yup, you got it. :)